OpenStudy (anonymous):

Soybean meal is 14% protein. cornmeal is 7% protein. How many of each should be mixed together to get 280lb. that is 9% protein?

OpenStudy (saifoo.khan):

Soybean meal is 14% protein; cornmeal is 7% protein. How many pounds of each should be mixed together to get 280-lb mixture that is 9% protein? ------------ Quantity Eq: S + C = 280 lb Protein Eq: 0.14S+0.07C = 0.09*280 ---------------------- Multiply thru 1st by 14 Multiply thru 2nd by 100 ----------------------- 14S + 14C = 14*280 14S + 7C = 9*280 --- Subtract and solve for "C": 7C = 5*280 C = 200 lb (amt. of cornmeal needed in the mix) --- Since S+C = 280, S = 80 lb (amt. of soybean needed in the mix)

OpenStudy (anonymous):

Awsome.. Thanks

OpenStudy (saifoo.khan):

Welks! :D